OpenStudy (anonymous):

In the following situations, is there likely to be a correlation? If so does the correlation reflect a causal relationship? Explain. the average daily winter temperature and your heating bill A. There is a positive correlation. The higher the average daily winter temperature the higher your heating bill. B. There is a negative correlation and a causal correlation. The higher the average daily winter temperature the lower your heating bill. C. There is no correlation.

OpenStudy (dumbcow):

B, you use less heat in warmer temperatures

OpenStudy (anonymous):

thanks what about Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10)

OpenStudy (dumbcow):

parallel --> same slope point slope form of line y - y1 = m(x-x1) plug in your slope (m) and given point (x1,y1) = (4,10)

OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6 A. y + 3 = 6(x - 8) B. y - 3 = 6(x - 8) C. y - 3 = 6(x + 8) D. y + 3 = 6x + 8

OpenStudy (anonymous):

which one

OpenStudy (dumbcow):

haha what happened to last question? anyway i gave you general form, you pick which one

OpenStudy (anonymous):

dnt get it

OpenStudy (dumbcow):

dont get what? plug in 8 and 3 for x1 and y1

OpenStudy (anonymous):

8/3

OpenStudy (dumbcow):

?

OpenStudy (dumbcow):

y - y1 = m(x-x1) m=6 x1 =8 y1 =3 plug those values in

OpenStudy (anonymous):

C. y - 3 = 6(x + 8)

OpenStudy (dumbcow):

how does (x-8) become (x+8) its x - x1

OpenStudy (dumbcow):

if x1 = -8, then it would be (x+8) x-(-8) = x+8